# GraphToxin: Reconstructing Full Unlearned Graphs from Graph Unlearning

### v1

@@ 4.2.2 Effectiveness
TABLE II: The Performance of GraphToxin for A Single Node Removal (Backbone: GCN). Unit: 1e-2. Arrow indicates the direction of better performance and the blue color denotes the superb ‘best’ results (in bold).
Dataset GU Baseline NRMSE($\downarrow$) ED($\downarrow$) PMGK($\uparrow$) ATT. ACC($\uparrow$) ATT. FID($\uparrow$) PWD($\downarrow$)
@@
4.2.3 Flexibility
To answer SQ3, we evaluate the performance of GraphToxin across different graph unlearning methods and GNN backbones.
Graph Unlearning From Table II, GraphToxin demonstrates superior and consistent performance across all unlearning methods. Specifically, in Cora, the attack against Retraining achieves the highest NRMSE and attack accuracy in both random and worst cases. It also excels in embedding distance, attack fidelity, and posterior Wasserstein distance (random case), and in PMGK and attack fidelity (worst case). The attack against MEGU outperforms others in embedding distance and posterior Wasserstein distance in the worst case, and in attack fidelity in the random case. In PubMed, GraphToxin produces very similar results in nearly all metrics, except for PMGK, where ETR stands out in the random case while Retraining leads in the worst case. In Photo, the attack against ETR performs better in NRMSE (both cases) and exceeds others in embedding distance and attack accuracy (worst case), and in PMGK (random case), while Retraining excels in PMGK, attack fidelity and posterior Wasserstein distance in the worst case. Despite these differences, the overall variations among unlearning methods are negligible, with only Retraining outperforming by a maximum of $2.5\%$ in attack accuracy in Cora.
GNN Backbones Since GraphToxin achieves consistent performance across different graph unlearning methods on three graph datasets, in this experiment, we directly assess the worst unlearning case—Retraining—on different GNN backbones. As shown in Table VI in Appendix .1.4, similarly, GraphToxin retains its superiority and generalization across all GNN backbones and graph datasets. In Cora, GraphToxin with the GCN backbone outperforms GraphSAGE and SGC in all performance-level metrics. It leads in NRMSE and PMGK in the random case and embedding distance in the worst case. GraphSAGE achieves the best NRMSE in the worst case and embedding distance in the random case, while SGC excels in PMGK in the worst case. In PubMed, GraphToxin with the GCN backbone performs better in embedding distance, PMGK and posterior Wasserstein distance (worst case), and surpasses others in attack accuracy (random case). GraphSAGE exceeds in NRMSE and attack fidelity (worst case), and in PMGK (random case), while SGC excels in NRMSE, embedding distance, attack fidelity and posterior Wasserstein distance (random case), and in attack accuracy (worst case). In Photo, GraphToxin with the GCN backbone achieves superior NRMSE and embedding distance in the random case, and in PMGK in both cases. GraphSAGE exceeds in attack accuracy, fidelity, and posterior Wasserstein distance in the random case, while SGC excels in NRMSE embedding distance, attack accuracy, fidelity, and posterior Wasserstein distance in the worst case.
In conclusion, there is no universally most vulnerable GNN due to the complex interplay among model architecture, graph sparsity and homophily. We provide the following plausible explanations: 1) GCN aggregates information over local neighborhoods. This paradigm creates a strong and predictable link between a node’s features and its neighbors for full graph recovery. However, this “imprint” may fail due to rapid aggregation expansion and subsequent over-smoothing. Hence, GCN obtains more competitive performance in small-scale, sparse and homophilous graphs like Cora. 2) Since GraphSAGE utilizes neighbor sampling rather than the full neighborhood aggregation, gradients are influenced by only a subset of the true local structure in each forward pass. This inherent randomness can obfuscate GraphToxin as it relies on incomplete gradients as guidance, especially in the sparse graphs. However, in dense graphs like Photo, GraphToxin with the GraphSAGE backbone can still achieve promising attack results. 3) SGC employs linear multi-hop feature propagation. While its linearity offers a direct and simpler proxy for full graph recovery, its linear smoothing can also destroy fine-grained and node-specific details. This dual effect explains why SGC stands out in some metrics in PubMed and Photo while its performance degrades in Cora.
@@ 4.3.1 Worst-case Analysis
In previous subsections, we addressed the necessity and importance of the worst-case evaluation. Here we shift the focus to its analysis. As shown in Table II and Table VI, interestingly, in Cora, GraphToxin against all graph unlearning methods achieves better performance in the worst case in nearly all metrics, except for PMGK. Conversely, in PubMed and Photo, GraphToxin performs betters in the random case in embedding distance, PMGK, attack fidelity and posterior Wasserstein distance. Additionally, GraphToxin with the GCN backbone also excels in attack accuracy in the random case.
For different GNN backbones, in Cora, GraphToxin in the random case achieves the best PMGK across all backbones. It also excels in embedding distance with SGC, and in attack accuracy, fidelity, and posterior Wasserstein distance with both GraphSAGE and SGC. In PubMed, it similarly leads in PMGK across all GNN backbones, and excels in embedding distance with SGC, in attack fidelity with GCN and SGC, and in posterior Wasserstein distance with SGC. In Photo, GraphToxin in the random case surpasses others in PMGK and posterior Wasserstein distance across all backbones and obtains the best embedding distance, attack accuracy, and fidelity with both GCN and GraphSAGE.
In the worst-case scenario, the top $10\%$ of nodes with the highest degrees are removed before GraphToxin attempts to reconstruct the full unlearned graph. Removing these highly influential structural hubs poses extra challenges, as the attack’s efficacy dependent on a delicate trade-off. While a moderate expansion of local neighborhoods can yield more informative gradients for graph recovery, these signals may also be distorted by over-smoothing or the loss of crucial semantic and structural details, particularly within heterogeneous neighborhoods or under linear smoothing effects.
4.3.2 Deletion Graph Size
Since we apply the same hyperparameter setting across all experiments to ensure a fair comparison, the reported results therefore represent a benchmark. We anticipate that case-specific hyperparameter tuning for each full recovered graph could yield further improvements in attack efficacy.
From Figure 5 in Appendix .1.5, we consistently observe that the attack performance is highly volatile when the unlearned graph size is small, particularly for embedding distance, attack accuracy and fidelity, and posterior Wasserstein distance. As the size increases, the attack performance tends to stabilize and converge across different GNN backbones. Specifically, GraphToxin with the GCN backbone excels in the majority of settings. While GraphSAGE and SGC can diverge from GCN when the graph size ranges from 203020\!-\!30 in Cora, their results ultimately converge as the graph size increases. NRMSE remains relatively flat and low, indicating its insensitivity to the deletion graph size. PMGK generally exhibits a negative correlation with the graph size. However, excluding GCN in Photo, its PMGK rebounds when the graph size becomes substantially large.
@@ 4.3.3 Node Homophily
From Figure 6 in Appendix .1.6, consistent attack performance is observed across diverse GNN backbones and graph datasets. Consistent with previous findings, GraphToxin with the GCN backbone achieves superior performance in nearly all metrics. However, GraphSAGE and SGC diverge from GCN in embedding distance, attack accuracy and fidelity, and posterior Wasserstein distance when node homophily approaches $1$. Outside of this extreme range, the attack performance across all GNN backbones tends to converge. We speculate that perfect or near-perfect homophilous neighborhoods may lead to non-distinct node embeddings, resulting in uninformative gradients that hinder full graph recovery. NRMSE remains consistently flat and low, indicating its general insensitivity to node homophily. PMGK, however, exhibits a slight positive correlation. Unlike Cora and Photo, the PMGK curve shows distinct turning points, suggesting a more nuanced, non-linear dependency on node homophily.
@@ 4.3.4 Fisher Coefficient
The curvature matching module is a key component in GraphToxin, hence, we compare GraphToxin with and without this module to analyze its contribution for full graph recovery. We vary its Fisher coefficient from $\{0,0.01,0.1,1,10\}$. Setting the coefficient to 0 is equivalent to an ablation study of the module. From Figure 3, the results empirically prove the effectiveness of curvature matching, as performance in most metrics improves when the coefficient is non-zero. Consistent with previous findings, targeting the worst-case node removal generally yields more effective performance than the random removal. However, there are several notable exceptions. The random case significantly outperforms the worst case in PMGK across all coefficients. Additionally, the random case excels in attack fidelity for the fisher coefficient $\alpha_{1}$ of $0.01$ and $0.1$, and in posterior Wasserstein distance when $\alpha_{1}\in\{0,0.01,0.1\}$. Overall, GraphToxin achieves competitive performance when $\alpha_{1}\in\{0.01,0.1,1\}$ and yields the relatively optimal results for both random and worst cases when $\alpha_{1}=0.1$.
@@ 4.3.5 Laplacian Coefficient
We further assess the contribution of the feature smoothing module. Since prior studies set their Laplacian coefficients ($\alpha_{2}$) to fairly small values [gradient_attack_gnn, graph_mia], we tune $\alpha_{2}\in\{0,1e\!-\!7,1e\!-\!5,1e\!-\!3,0.1\}$. From Figure 4, the attack results demonstrate the effectiveness of this module. We also observe that targeting the worst-case node removal outperforms the random removal in nearly all metrics. The random case only surpasses the worst case in PMGK and attack accuracy, fidelity, and posterior Wasserstein distance when $\alpha_{2}=0$. Overall, GraphToxin achieves superior performance when $\alpha_{2}\in\{1e\!-\!7,1e\!-\!5\}$.
@@ 4.4 Extensions (SQ6 and SQ7)
Due to space limitations, we provide detailed performance and analysis of the extensions to multiple node removals and the black-box setting in Appendix .1.7 and .1.8, respectively.
@@ 4.5.1 Defense (SQ8)
To investigate the performance of various defense mechanisms against GraphToxin, we utilize the most vulnerable and stable attack target—Retraining after a single node removal in the white-box setting. We adopt both in-processing and post-processing privacy-preserving techniques. Pre-processing techniques are out of our scope as they might cause obfuscation during node removals.
@@
Node-level Differential Privacy (Node-DP) in GNNs Node-DP in GNNs provides a formal privacy guarantee that protects an individual node and all its incident edges during message passing. Achieving this guarantee is substantially more difficult than enforcing standard DP or edge-level DP, as nodes introduce structural coupling and multi-hop dependencies that amplify per-sample sensitivity and complicate privacy accounting. Currently, there is no unified node-DP framework adapted to diverse GNN backbones. Existing studies either inject noises to shallow GNNs [1_layer_dp] or decouple graph convolutions through different architectures [gap, dpar]. Furthermore, a larger privacy budget is typically consumed for a better privacy and utility trade-off [dp_amplify, personal_node_dp, node_dp]. Consequently, realizing practical node-DP with competitive utility still remains an active research frontier.
@@
Following Xiang et al. [node_dp], we adopt HeterPoison to achieve node-DP. As their default privacy setting results in unacceptable utility in our GNN backbones, we choose to tune the $\sigma\in[0.006,0.007,0.008,0.009,0.01]$ in the symmetric multivariate Laplace noise $\mathcal{LAP}(0,\sigma^{2}\mathbb{I}^{d})$. This tuning yields the test accuracy of 79.40%, 76.00%, 69.90%, 61.00% and 51.90%, respectively. To balance utility and privacy, we select $\sigma=0.007$ to fight against GraphToxin.
From Table III, even a rigorous privacy protection against GNNs cannot sufficiently defend GraphToxin. Under the node-DP guarantee, GraphToxin still retains competitive performance compared to the pure attack without any defense. Surprisingly, the attack performance even improves in attack accuracy. We speculate that this phenomenon occurs because DP noises can preferentially disrupt weak and unstable signals within the gradients, thus forcing GraphToxin to recover more robust graph components from more informative and stable gradients. While these components may not form a perfect replica of the original graph, they can contribute to the final node classification. In this sense, node-DP may act as a form of regularization, preventing GraphToxin from overfitting to trivial details.
@@
Gradient Compression (GC). Given that GraphToxin exploits the real or estimated gradient difference for full graph recovery, it is imperative to enhance the privacy of released gradients. To this end, gradient compression techniques are widely employed against such gradient-based attacks [gia_survey]. Following Vogels et al. [powersgd], we prune the smaller values to 0 by a $p$ proportion to assess the impact of gradient sparsity on GraphToxin. We experiment with high pruning ratios of $p\in[0.5,0.7,0.9]$. The most stringent case, $p=0.9$, enforces the extreme gradient sparsity where only very few gradients are preserved. The defense results are shown in Table III. Similarly to node-DP, gradient compression cannot effectively defend GraphToxin. Interestingly, preserving only a fraction of the gradients can lead to comparable or even superior attack performance compared to using the full gradients as a proxy. When the pruning ratio increases, the overall attack performance slightly declines and then improves. Even in the most stringent scenario, GraphToxin under this privacy protection technique can exceed that of the undefended attack. This specific finding further reinforces our previous speculation: only retaining important and informative gradients rather than replicating trivial details can enhance the attack performance of GraphToxin.
@@
TABLE III: The Performance of Node-DP and Gradient Compression in Defending GraphToxin. Unit: 1e-2. Arrow indicates the direction of better performance and the bold font denotes the ‘best’ results.
Defense Type NRMSE($\downarrow$) ED($\downarrow$) PMGK($\uparrow$) ACC($\uparrow$) FID($\uparrow$) PWD($\downarrow$)
@@ 4.5.2 Limitation and Future Work
Assumption Relaxation Similar to existing studies [gradient_attack_gnn, first_unlearn_attack], we assume that attackers can master node labels and the size of unlearned samples. Despite its feasibility, it is beneficial to propose a more advanced GU-FGRA without such information. A possible solution is to integrate label inference attacks into the GU-FGRA framework. When the size of the unlearned graphs is unknown, a Frobenious norm regularizer [gradient_attack_gnn, graph_mia] can be introduced to ensure graph sparsity. Furthermore, graph condensation-related metrics can be utilized to measure attack performance as the ground-truth deletion graph can be treated as an approximate condensed version of the recovered graph. We leave this investigation as future work.
Large-scale Node Removal This study mainly focuses on the effectiveness of GraphToxin in the context of small-scale multiple node removals. Although we explore the challenging worst-case scenario by simultaneously removing multiple highest-degree nodes, which results in a full deletion graph with a relatively large size, investigating the scalability of GraphToxin to large-scale multiple node removals presents a crucial direction for future work.
More Worst-case Analysis Since the degree of deleted nodes directly affects graph unlearning performance, we remove nodes with the highest degrees as the worst case in this study. Other metrics such as node influence [node_influence, GIF] can also be utilized for the worst-case analysis.
Deep Investigation on Complex Interplays We identify several interesting and counter-intuitive findings in this research. For instance, GraphToxin in the worst-case scenario achieves better overall attack performance than the random case. Furthermore, GraphToxin under a rigorous DP guarantee still retains competitive or even outstanding attack performance, and only preserving an extremely small fraction of the gradients can yield comparable or even superior attack performance compared to using the full gradients. An in-depth investigation should be conducted due to the complex interplays among model architectures, graph sparsity, node homophily and other possible factors.

### v2

@@ 4.2.2 Effectiveness (SQ2)
Table 2: The Performance of GraphToxin for single-node Removal (Backbone: GCN). Unit: 1e-2. Arrow indicates the direction of better performance and the blue color denotes the superb ‘best’ results (in bold).
Dataset GU Baseline RNMSE($\downarrow$) ED($\downarrow$) PMGK($\uparrow$) ATT. ACC($\uparrow$) ATT. FID($\uparrow$) PWD($\downarrow$)
@@
4.2.3 Flexibility (SQ3)
We evaluate the flexibility of GraphToxin across different graph unlearning methods and GNN backbones.
Graph Unlearning The results in Table 2 demonstrate the stable and consistent performance of GraphToxin across all graph unlearning methods. While minor variations are observed, these differences arise from the interaction between how unlearning signals are manifested in the parameter space and the structural properties of the underlying graphs. Specifically, the attack against Retraining tends to induce larger parameter changes, which are particularly exploitable in small and sparse graphs such as Cora. In contrast, MEGU introduces learning-based modules to suppress large parameter deviations. This design mitigates the propagation of unlearning signals and can lead to reduced distinguishability under structure-sensitive metrics such as PMGK. ETR adopts a training-free and highly localized strategy that erases the influence of deleted nodes while rectifying collateral damage to neighboring nodes. Therefore, ETR can better preserve local structures, which may account for its stronger global-level alignment across diverse graphs.
GNN Backbones We next fix the unlearning target as Retraining and investigate the impact of different GNN backbones. Compared to the nearly identical performance observed across unlearning methods in Table 2, GraphToxin exhibits distinct performance across GNN backbones, while still maintaining its overall superiority and generalization across all graph datasets (Table 6). Detailed results and analysis are provided in Appendix B.4 due to space constraints.
@@ 4.3.1 Worst-case Analysis
Previous subsections motivate the necessity of worst-case evaluation. We now analyze its impact on attack performance. As summarized in Table 2 and Table 6, worst-case node removal leads to markedly different behaviors across graphs. On Cora, GraphToxin consistently achieves stronger performance under worst-case removal across nearly all metrics, except PMGK. In contrast, on PubMed and Photo, GraphToxin generally performs better under random-node removal for global and performance-level metrics. Similar trends are observed across different GNN backbones.
This divergence highlights the inherent trade-off in worst-case evaluation. Removing high-degree nodes increases reconstruction difficulty by amplifying structural ambiguity and semantic misalignment. While moderate neighborhood expansion can yield more informative signals for recovery, these signals may be distorted by over-smoothing or losing crucial semantic and structural details during information propagation, particularly within heterogeneous neighborhoods.
Figure 3: Impact of the Full Unlearned Graph Size and Node Homophily
4.3.2 Unlearned Graph Size
Figure 3 shows that the attack performance is unstable when the unlearned graph is small. As the graph size increases, the performance stabilizes and gradually converges across different GNN backbones. We provide the visualization of rest metrics and corresponding analysis in Appendix B.5.
@@ 4.3.3 Node Homophily
As illustrated in Figure 3, GraphToxin demonstrates stable performance under varying homophily levels, with the GCN backbone achieving the strongest performance across most metrics. When node homophily approaches one, GraphSAGE and SGC diverge from GCN mainly in performance-level evaluation. Outside this extreme range, the attack performance across all backbones tends to converge. One possible explanation is that near-perfect homophilic neighborhoods produce indistinguishable node embeddings, thus leading to noninformative gradients for recovery. Similarly, the visualization and analysis of rest metrics are presented in Appendix B.6.
@@ 4.3.4 Fisher Coefficient
We evaluate the contribution of the curvature matching module by varying its Fisher coefficient ($\alpha_{1}$) from $\{0,0.01,0.1,1,10\}$, where $\alpha_{1}=0$ serves as the ablation. Figure 4 empirically confirms the effectiveness of curvature matching, as the attack performance generally improves when the coefficient is non-zero. Consistent with previous findings, the results under worst-case removal typically exceed those under random removal. Additionally, the random case achieves higher attack fidelity when $\alpha_{1}\in\{0.01,0.1\}$ and lower posterior Wasserstein distance when $\alpha_{1}\in\{0,0.01,0.1\}$. Overall, GraphToxin attains competitive performance when $\alpha_{1}\in\{0.01,0.1,1\}$ and yields better results for both random and worst cases when $\alpha_{1}=0.1$.
@@ 4.3.5 Laplacian Coefficient
Figure 5 demonstrates the effectiveness of the feature smoothing module. Similar to prior studies that set Laplacian coefficients ($\alpha_{2}$) to fairly small values [gradient_attack_gnn, graph_mia], we tune $\alpha_{2}\in\{0,1e\!-\!7,1e\!-\!5,1e\!-\!3,0.1\}$. We observe that the attack performance under the worst-case removal exceeds that under the random removal in nearly all metrics while the random case achieves stronger performance mainly in global and performance-level metrics. Overall, GraphToxin achieves superior performance when $\alpha_{2}\in\{1e\!-\!7,1e\!-\!5\}$.
@@ 4.4 Extensions (SQ6 and SQ7)
Due to space limitations, we provide detailed performance and analysis of the extensions to multiple-node removal and the black-box setting in Appendix B.7 and B.8, respectively.
@@ 4.5.1 Defense (SQ8)
To investigate the performance of defense mechanisms against GraphToxin, we use the most vulnerable and stable attack target—Retraining after single-node removal in the white-box setting. We adopt both in-processing and post-processing privacy-preserving techniques. Pre-processing techniques are out of our scope as they typically alter graphs before training, making it difficult to disentangle whether observed effects stem from unlearning or from data modification.
@@
Node-level Differential Privacy (Node-DP) in GNNs Node-DP in GNNs aims to protect an individual node and all its incident edges during message passing. Compared to standard or edge-DP, enforcing node-DP is substantially more challenging due to structural coupling and multi-hop dependencies, which amplify per-sample sensitivity and complicate privacy accounting. Existing studies either inject noises to shallow GNNs [1_layer_dp] or redesign architectures to decouple graph convolutions [gap, dpar]. Despite these efforts, achieving strong node-level privacy guarantees with competitive utility remains an open problem [dp_amplify, personal_node_dp, node_dp].
@@
Following Xiang et al. [node_dp], we adopt HeterPoison to implement node-DP. As the default configuration leads to severe utility degradation on GNN backbones, we directly tune the noise scale $\sigma\in\{0.006,0.007,0.008,0.009,0.01\}$ from the symmetric multivariate Laplace noise $\mathcal{LAP}(0,\sigma^{2}\mathbb{I}^{d})$, resulting in test accuracies of 79.40%, 76.00%, 69.90%, 61.00% and 51.90%, respectively. We select $\sigma=0.007$ to balance between privacy and utility.
From Table 3, node-DP fails to effectively mitigate GraphToxin. The reconstruction quality remains comparable to the undefended setting, and attack accuracy may even increase. We attribute this behavior to the injected DP noise, which suppresses weak and unstable gradient components while preserving more informative signals for recovery. In this sense, node-DP can act as a form of regularization, preventing GraphToxin from overfitting to trivial details.
@@
Gradient Compression (GC). Given that GraphToxin exploits the real or surrogate gradient difference for recovery, it is imperative to enhance the privacy of released gradients. Gradient compression techniques are widely adopted against such gradient-based attacks [gia_survey]. Following Vogels et al. [powersgd], we prune a proportion $\mathrm{p}$ of smaller gradients to assess the impact of gradient sparsity on GraphToxin. We experiment with high pruning ratios of $\mathrm{p}\in\{0.5,0.7,0.9\}$. The most stringent case, $\mathrm{p}=0.9$, enforces the extreme gradient sparsity where only very few gradients are preserved. The results are shown in Table 3. Similar to node-DP, GC cannot effectively defend GraphToxin. Preserving only partial gradients can lead to attack performance comparable to or even exceeding that obtained from full gradients. Even under extreme sparsification, GraphToxin remains competitive relative to the undefended baseline. This finding further supports our previous speculation that retaining only important and informative gradients rather than replicating trivial details can enhance the attack performance of GraphToxin.
@@
Table 3: The Performance of Node-DP and Gradient Compression in Defending GraphToxin (Dataset: Cora). Unit: 1e-2. Arrow indicates the direction of better performance and the bold font denotes the ‘best’ results.
Defense Type RNMSE($\downarrow$) ED($\downarrow$) PMGK($\uparrow$) ACC($\uparrow$) FID($\uparrow$) PWD($\downarrow$)
@@ 4.5.2 Limitation and Future Work
Assumption Relaxation Consistent with prior work [gradient_attack_gnn, first_unlearn_attack], we assume that the attacker has access to node labels and the size of the unlearned subgraph. While this assumption is standard and feasible, removing it would further enhance the applicability of GU-FGRA. One possible direction is to incorporate label inference into the GU-FGRA pipeline. When the size of the full unlearned graph is unknown, a Frobenious norm regularizer [gradient_attack_gnn, graph_mia] can be introduced to ensure graph sparsity. Furthermore, graph condensation metrics can be utilized to evaluate reconstruction quality as the ground-truth unlearned graph can be treated as an approximate condensed target. We leave these extensions for future work.
Large-scale Node Removal This work follows existing graph unlearning research and primarily examines the effectiveness of GraphToxin under relatively small-scale node removal. Although we explore the challenging worst-case scenarios that generate comparatively larger unlearned graphs by simultaneously removing multiple highest-degree nodes, investigating the scalability of GraphToxin under large-scale multiple-node removal presents a promising direction for future work.
Beyond Degree-Based Worst-case Analysis Node degree provides a direct and interpretable proxy for structural influence in the worst-case evaluation, as deleting a node necessarily removes all of its incident edges. This correspondence makes degree a natural criterion for modeling the immediate structural impact of unlearning requests, while other measures such as node influence [node_influence, GIF], may capture different worst-case behaviors. Exploring alternative influence criteria could provide a more comprehensive view of GU-FGRA.
Deep Investigation on Complex Interactions Several non-trivial findings emerge from the experimental results. For instance, GraphToxin remains effective under strict differential privacy guarantees, and retaining only a small fraction of gradients can suffice for GU-FGRA. These observations suggest complex interactions among model architectures, graph sparsity, node homophily and other possible factors. A deeper investigation on these factors is left for future work.